# Dependence of the critical temperature and disorder in holographic superconductors on superfluid density

### v3

@@ 3.1 Phase boundaries
As we have mentioned, we have two choices of the normalizable modes. We consider the following ansatz of the perturbation for each theory. For the $O_{1}$ theory, we write $\Phi(z)$ as
$\Phi(z)=\frac{\langle O_{1}\rangle}{\sqrt{2}}\frac{z}{r_{h}}F(z),\quad F(0)=1,\quad F^{\prime}(0)=0.$ (22)
For the $O_{2}$ theory, we write $\Phi(z)$ as
$\Phi(z)=\frac{\langle O_{2}\rangle}{\sqrt{2}}\frac{z}{r_{h}^{2}}F(z),\quad F(0)=0,\quad F^{\prime}(0)=1.$ (23)
We also impose the regular condition for $F(z)$ at $z=1$. The equation of motion (17) becomes
$F^{\prime\prime}(z)+\left(\frac{4}{z}+\frac{f^{\prime}(z)}{f(z)}+\frac{h^{\prime}(z)}{h(z)}\right)F^{\prime}(z)+\\ \left(\frac{2}{z^{2}}+\frac{h^{\prime}(z)}{zh(z)}+\frac{r_{h}^{2}\left(q^{2}A_{t}(z)^{2}-B(\phi(z))f(z)\right)}{z^{4}f(z)^{2}}+\frac{f^{\prime}(z)}{zf(z)}\right)F(z)=0\ .$ (24)
This can be understood as the Strum-Liouville (SL) problem. The equation in the SL form becomes
@@
$\begin{split}p(z)&=z^{4}f(z)h(z),\\ q(z)&=z^{3}f(z)h(z)\left(\frac{2}{z}+\frac{f^{\prime}(z)}{f(z)}+\frac{h^{\prime}(z)}{h(z)}\right)-r_{h}^{2}B(\phi(z))h(z),\\ r(z)&=\frac{(-1+z)^{2}r_{h}^{4}h(z)}{f(z)},\qquad\lambda=\frac{q^{2}\mu^{2}}{r_{h}^{2}}.\end{split}$ (26)
We can solve the above equation numerically. The minimum eigenvalue $\lambda$ will correspond to the critical value of $\mu$ or $T_{c}$ after rescaling parameters. Moreover, we can also study the approximation for the lowest eigenvalue, as follows. Supposing that $\Psi_{n}$ is the eigenfunction for $n$-th eigenvalue $\lambda=\lambda_{n}$, we can formally write $\lambda_{n}$ as
@@
This is called Rayleigh Quotient. We leverage several properties of the Strum-Liouville eigenvalue problem. One of them is the set of eigenfunctions is complete, i.e., $F(z)\sim\sum_{n=1}^{\infty}c_{n}\Psi_{n}(z)$. Thus, we can generalize this formula for a trial function $F(z)$ which is not a solution of the ordinary differential equation (25) but satisfies the boundary conditions. In this study, we consider trial functions with one parameter $\alpha$ denoted by $F_{\alpha}$. In the SL problem, the smallest eigenvalue $\lambda_{1}$ always exists. Correspondingly, $\lambda_{\alpha}$, associated with the trial function $F_{\alpha}$, also has a minimum. We write
@@
This becomes an estimation for the actual smallest eigenvalue: $\hat{\lambda}_{1}\approx\lambda_{1}$.
We have to note that the background spacetime also exhibits the phase transition between Eq. (5) and Eq. (7). Since we are working in the probe limit, this phase transition occurs at $T=k/(2\sqrt{2}\pi)$, regardless the charged-scalar instability. In the following, we first explore the charged-scalar instability in each background geometry. After that, we will show the correct phase boundaries as Fig. 3 by combining the results obtained in each geometry.
@@
It satisfies the boundary conditions, $F(0)=1$ and $F^{\prime}(0)=0$. We obtain the approximate formula for the minimal eigenvalue $\hat{\lambda}_{1}$. Figure 1 shows the phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane obtained by the approximation, and also by the numerical way for comparison. Note that $q$ can be scaled out in the probe limit analysis.222 In the backreacted case, $q$ can no longer be scaled out; different $q$ explains different setups. Similar to the original model of the holographic superconductor, it is expected that the probe limit corresponds to $q\to\infty$ limit. See Appendix A. The approximation agrees with the numerical result well. Since the full expression of the approximation is too complicated, we do not exhibit it here. In the $O_{1}$ theory, the phase boundaries always lie in on $T>k/(2\sqrt{2}\pi)$ in each background geometry, as shown in Fig. 1. As a result, only the result in the bald black hole is relevant because the bald black hole is the ground state of the gravity sector when $T>k/(2\sqrt{2}\pi)$.
Refer to captionRefer to caption
@@
Figure 1: Superconducting phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane for the $O_{1}$ theory in the two different black hole solutions. The black dotted lines show $T=k/(2\sqrt{2}\pi)$ where the phase transition of the background spacetime occurs. (left) The result in the bald black hole (7). (right) the result in the dilatonic black hole (5). The curves show the SL approximations while the points show the direct numerical results. In the right panel, the results depend on $\tau$. The curves and data correspond to $\tau=0,\frac{5/4}{\sqrt{3}},\frac{1}{\sqrt{3}}$ from bottom to top.
Refer to captionRefer to caption
@@
Figure 2: Superconducting phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane for $O_{2}$-theory in the two different black hole solutions. The black dotted lines show $T=k/(2\sqrt{2}\pi)$ where the phase transition of the background spacetime occurs. (left) The result in the bald black hole (7). (right) The result in the dilatonic black hole (5). The curves show the SL approximations, while the points show numerical results obtained directly. In the right panel, the results depend on $\tau$. The curves and data correspond to $\tau=0,\frac{3}{4\sqrt{3}}$ and $\frac{1}{\sqrt{3}}$ from bottom to top.
@@
In the case of the bald black hole, the curve is parameterized by $k/r_{h}$ in $0<k/r_{h}<\sqrt{2}$. In $k/r_{h}\to\sqrt{2}$, the curve approaches to the dotted line denoting $T=k/(2\sqrt{2}\pi)$. In the case of the dilatonic black hole, the curves are parameterized by $P/r_{h}$ in $-\infty<P/r_{h}<0$, and small $|P|/r_{h}$ corresponds to large $\frac{k}{q\mu}$. Expanding the approximation for $\tau=0$ in small $\epsilon\equiv-P/r_{h}$, we obtain
$\lambda_{F_{\alpha}}=\frac{1}{4\ln 2-2}\epsilon+\order{\epsilon^{2}}.$ ) . (30)
@@
$T_{c}$ approaches to the critical value of the phase transition of the background spacetime. In the case of the bald black hole, $T_{c}$ approaches to this value for large $k$, too. It can be also confirmed from the numerical results for $\frac{T_{c}}{q\mu}$ at large $\frac{k}{q\mu}$ (, see Fig. 1). This result illustrates the critical temperature is directly proportional to the momentum dissipation strength in the $O_{1}$ theory. Comparable outcomes were indeed numerically established in the studies of [54] and [46], yet these works did not explicitly elucidate or demonstrate such findings within the regime transcending the probe limit (refer to Figure 1 presented in both [54] and [46] for a clearer understanding).
@@
$F_{\alpha}(z)=z\left(1+\alpha z\right)\left(1-\frac{Pz}{r_{h}}\right)^{-\frac{\sqrt{3}}{2}\tau}.$ (34)
@@
This trial function satisfies $F(0)=0$ and $F^{\prime}(0)=1$ so it is suitable for studying the $O_{2}$ theory.333 One can also consider $F_{\alpha}=\frac{z\left(1+\alpha z\right)}{1+z}\left(1-\frac{Pz}{r_{h}}\right)^{-\frac{\sqrt{3}}{2}\tau}$, which gives good agreement too. The exponential factor $-\frac{\sqrt{3}}{2}\tau$ is just same as those involved in the coupling term $B(\phi)$, see Eq. (11). Figure 2 shows the results of the phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane obtained by the approximation and the numerical method, for a comparison. The approximation almost agrees with the numerical result. Unlike the $O_{1}$ theory, the curves cross the line of $T=k/(2\sqrt{2}\pi)$, and $T_{c}$ reaches zero if $\tau$ is small. In the case of the dilatonic black hole, the curves are parameterized by $P/r_{h}$, but it can take a value in $10<P/r_{h}<$ 1 in the $O_{2}$ theory. At $P/r_{h}=0$, the result does not depend on $\tau$ which is located at $(\frac{k}{q\mu},\frac{T_{c}}{q\mu})\approx(0.41,0.049)$ on the dotted line. For large enough $\tau$, $T_{c}$ linearly depends on $k$ but the coefficient is different from Eq. (32). For $\tau=1/\sqrt{3}$, we numerically obtain $T_{c}\approx 0.0193k$ for large $k$. Such behavior in $T<k/(2\sqrt{2}\pi)$ is qualitatively same as those studied in Ref. [46].
@@
Finally, we show the correct phase boundaries for each theory. Combining the results in Fig. 1 for the $O_{1}$ theory, and Fig. 2 for the $O_{2}$ theory, we obtain the phase boundaries as Fig. 3. According to the critical value of $T=k/(2\sqrt{2}\pi)$, which is shown as the dotted line, we have switched the using background geometry and the corresponding results. Note that the purple curves in $T>k/(2\sqrt{2}\pi)$ are independent of the choice of $\tau$ because $\tau$ represents a coupling with the dilaton. For the $O_{1}$ theory, we employed only the purple curve corresponding to the bald black hole because all curves are located in $T>k/(2\sqrt{2}\pi)$. The result for the $O_{2}$ theory corresponds to the fully backreacted result in Ref. [46]. We also provide further comparisons with the full analysis in Appendix A. The calculation of the critical temperature $T_{c}$ in this study demonstrates that employing the probe limit yields coherent and meaningful results by duly considering the phase transition dynamics within the background gravity sector.
@@
For both the $O_{1}$ theory and the $O_{2}$ theory with substantial $\tau$ values, it is revealed that $T_{c}$ increases as the parameter $k$ grows at high $k$ regimes. Specifically, for the $O_{1}$ theory, the large-$k$ behavior can be effectively approximated using Eq. (32). Although initially, this positive correlation between $T_{c}$ and $k$, which signifies the strength of momentum dissipation, may seem counter-intuitive, such a trend has been consistently observed across various holographic models incorporating $k$ [46, 54], as previously discussed. In light of these findings, a future discourse on the validity and implications of this scaling relationship between $T_{c}$ and $k$ from the perspective of condensed matter physics would indeed be valuable and insightful. It remains to be explored whether this phenomenon aligns with physical expectations or presents novel challenges to our current understanding.
Refer to captionRefer to caption
@@ 3.2 Condensation
Figure 4 shows the condensates $\expectationvalue{O_{i}}$ as functions of $T$ for $O_{i}$ theories. We set $\tau=0$ for simplicity. To obtain this figure, we have changed the using background geometry at the phase transition point $T=k/(2\sqrt{2}\pi)$. We show this point as a small circle on the curve, if it exists below the superconducting phase transition point. The condensate $\expectationvalue{O_{1}}$ grows as temperature decreases, whereas $\expectationvalue{O_{2}}$ remains finite at $T\to 0$. In particular, there is a lower bound for the possible $T$ in the $O_{1}$ theory. This behavior is same as those studied in [23], and is a limitation in the probe limit analysis.444 In the $O_{2}$ theory with $\tau=0$, $T\to 0$ limit is possible. Since the lowest temperature is always larger than $T=k/(2\sqrt{2}\pi)$ for the $O_{1}$ theory, the results in the bald black hole, denoted as solid curves, are favored. We expect that the low-temperature divergence of $\expectationvalue{O_{1}}$ will be cured by considering beyond the probe limit, similarly to Ref. [24]. One can see that $\expectationvalue{O_{1}}$ is enhanced by increasing $k$ for fixed $\mu$. On the other hand, $\expectationvalue{O_{2}}$ is suppressed as $k$ increases. These behaviors of the condensates are consistent with the relation between $T_{c}$ and $k$ for each theory, shown as Fig. 3.
Refer to captionRefer to caption

### v4

@@ 3.1 Phase boundaries
As we have mentioned, we have two choices of the normalizable modes. We consider the following boundary conditions for each theory. For the $O_{1}$ theory, we write $\Phi(z)$ as
$\Phi(z)=\frac{\langle O_{1}\rangle}{\sqrt{2}}\frac{z}{r_{h}}F(z),\quad F^{\prime}(0)=0,$ (22)
and the normalization is fixed by $F(0)=1$. For the $O_{2}$ theory, we write $\Phi(z)$ as
$\Phi(z)=\frac{\langle O_{2}\rangle}{\sqrt{2}}\frac{z}{r_{h}^{2}}F(z),\quad F(0)=0,$ (23)
and the normalization is fixed by $F^{\prime}(0)=1$. We also impose the regular condition for $F(z)$ at $z=1$. The problem can be understood as a Sturm-Liouville (SL) problem. With Eq. (20), the equation of motion (17) becomes
@@
$\begin{split}p(z)&=z^{4}\tilde{f}(z)\tilde{h}(z),\\ q(z)&=z^{3}\tilde{f}(z)\tilde{h}(z)\left(\frac{2}{z}+\frac{\tilde{f}^{\prime}(z)}{\tilde{f}(z)}+\frac{\tilde{h}^{\prime}(z)}{\tilde{h}(z)}\right)-B(\phi)\tilde{h}(z),\\ r(z)&=\frac{(1-z)^{2}\tilde{h}(z)}{\tilde{f}(z)},\qquad\lambda=\frac{q^{2}\mu^{2}}{r_{h}^{2}}.\end{split}$ (25)
We also defined $\tilde{f}(z)=r_{h}^{-2}f(z)$, and $\tilde{h}(z)=r_{h}^{-2}h(z)$.
In order to obtain $T_{c}$, we need finding solution that is regular at the horizon, and satisfies Eq. (22) or (23) at the AdS boundary, corresponding to the $O_{1}$ or $O_{2}$ theory. This is an eigenvalue problem with an eigenvalue $\lambda$. It can be solved numerically by using the shooting method. In the case of fixing $\mu$, i.e., grand canonical ensemble, we write
$\frac{T_{c}}{q\mu}=\begin{cases}\frac{1}{2\pi}\sqrt{\frac{1-P/r_{h}}{\lambda_{c}}}&\text{(dilatonic)}\\ \frac{6-k^{2}/r_{h}^{2}}{8\pi\sqrt{\lambda_{c}}}&\text{(bald)}\end{cases},\quad\frac{k}{q\mu}=\begin{cases}\frac{\sqrt{2}}{\sqrt{\lambda_{c}}}&\text{(dilatonic)}\\ \frac{k/r_{h}}{\sqrt{\lambda_{c}}}&\text{(bald)}\end{cases},$ (26)
by using (6) and (8). The largest $T_{c}$ is usually related to the phase transition, so we find the smallest $\lambda_{c}$, correspondingly.
Moreover, we can also study the approximation for the lowest eigenvalue, as follows. A similar method was utilized in Siopsis:2010uq . Supposing that $F=\Psi_{n}$ is the eigenfunction for $n$-th eigenvalue $\lambda=\lambda_{n}$, we can formally write $\lambda_{n}$ as
@@
We leverage several properties of the Sturm-Liouville eigenvalue problem. One of them is the set of eigenfunctions is complete. Thus, any regular function satisfying the boundary conditions can be expanded as $\sum_{n=1}^{\infty}c_{n}\Psi_{n}(z)$, where $c_{n}$ are coefficients of the expansion. We can generalize this formula (27) for a trial function, which is not an exact solution of the ordinary differential equation (25) but satisfies the boundary conditions. In this study, we consider trial functions with one parameter $\alpha$ denoted by $F_{\alpha}(z)$. We also define
@@
The function $\lambda_{(\alpha)}$ is bounded from below corresponding to that there is a minimum eigenvalue for the SL problem. We can find the minimum value $\hat{\lambda}_{c}$, and it becomes an estimation for the actual smallest eigenvalue $\lambda_{c}$.
We remark that the background spacetime also exhibits the phase transition between Eq. (5) and Eq. (7). Since we are working in the probe limit, this phase transition occurs at $T=k/(2\sqrt{2}\pi)$, regardless the charged-scalar instability. In the following, we first explore the charged-scalar instability in each background geometry. After that, we will show the correct phase boundaries as Fig. 3 by combining the results obtained in each geometry.
@@
It satisfies the boundary conditions, $F(0)=1$ and $F^{\prime}(0)=0$. Substituting this into (28), we obtain the approximate formula for the minimal eigenvalue, $\hat{\lambda}_{c}$. Subsequently, we obtain $\frac{T_{c}}{q\mu}$ and $\frac{k}{q\mu}$ by using Eq. (26) with $\hat{\lambda}_{c}$. The results are parameterized by $k/r_{h}$ in the bald black hole, and $P/r_{h}$ in the dilatonic black hole, respectively. We illustrate phase boundaries by plotting these one-parameter curves in each background geometry. Figure 1 shows the phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane obtained by the approximation, and also by the numerical way for comparison. Note that $q$ can be scaled out in the probe limit analysis.222 In the backreacted case, $q$ can no longer be scaled out; different $q$ explains different setups. It is expected that the probe limit corresponds to $q\to\infty$ limit. See Appendix A. The approximation agrees with the numerical result well. In the $O_{1}$ theory, the phase boundaries always lie in on $T>k/(2\sqrt{2}\pi)$ in two background geometries, as shown in Fig. 1. As a result, only the result in the bald black hole is relevant because the bald black hole is the ground state of the gravity sector when $T>k/(2\sqrt{2}\pi)$.
@@
Figure 1: Superconducting phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane for the $O_{1}$ theory in the two different black hole solutions. Each gray region indicates the thermodynamically unstable regime of the each background spacetime; the border is given by $T=k/(2\sqrt{2}\pi)$. (left) The result in the bald black hole (7). (right) the result in the dilatonic black hole (5). The curves show the SL approximations while the points show the direct numerical results. In the right panel, the results depend on $\tau$. The curves and data correspond to $\tau=0,\frac{3/4}{\sqrt{3}},\frac{1}{\sqrt{3}}$ from bottom to top.
@@
Figure 2: Superconducting phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane for $O_{2}$-theory in the two different black hole solutions. Each gray region indicates the thermodynamically unstable regime of the each background spacetime; the border is given by $T=k/(2\sqrt{2}\pi)$. (left) The result in the bald black hole (7). (right) The result in the dilatonic black hole (5). The curves show the SL approximations, while the points show numerical results obtained directly. In the right panel, the results depend on $\tau$. The curves and data correspond to $\tau=0,\frac{3}{4\sqrt{3}}$ and $\frac{1}{\sqrt{3}}$ from bottom to top.
@@
In the case of the bald black hole, the curve is parameterized by $k/r_{h}$ in $0<k/r_{h}<\sqrt{2}$. In $k/r_{h}\to\sqrt{2}$, the curve approaches to $T=k/(2\sqrt{2}\pi)$. In the case of the dilatonic black hole, the curves are parameterized by $P/r_{h}$ in $-\infty<P/r_{h}<0$, and small $|P|/r_{h}$ corresponds to large $\frac{k}{q\mu}$. Expanding the approximation for $\tau=0$ in small $\epsilon\equiv-P/r_{h}$, we obtain
$\hat{\lambda}_{c}=\frac{1}{4\ln 2-2}\epsilon+\order{\epsilon^{2}}.$ ) . (30)
@@
$T_{c}$ approaches to the critical value of the phase transition of the background spacetime. In the case of the bald black hole, $T_{c}$ approaches to this value for large $k$, too. It can be also confirmed from the numerical results for $\frac{T_{c}}{q\mu}$ at large $\frac{k}{q\mu}$ (see Fig. 1). This result illustrates the critical temperature is directly proportional to the momentum dissipation strength in the $O_{1}$ theory.
@@
[…]
@@
This trial function satisfies $F(0)=0$ and $F^{\prime}(0)=1$ so it is suitable for studying the $O_{2}$ theory.333 One can also consider $F_{\alpha}=\frac{z\left(1+\alpha z\right)}{1+z}\left(1-\frac{Pz}{r_{h}}\right)^{-\frac{\sqrt{3}}{2}\tau}$, which gives good agreement too. The exponential factor $-\frac{\sqrt{3}}{2}\tau$ is just same as those involved in the coupling term $B(\phi)$, see Eq. (11). Figure 2 shows the results of the phase boundaries in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane obtained by the approximation and the numerical method, for a comparison. The approximation almost agrees with the numerical result. Unlike the $O_{1}$ theory, the curves cross the line of $T=k/(2\sqrt{2}\pi)$, and $T_{c}$ reaches zero if $\tau$ is small. In the case of the dilatonic black hole, the curves are parameterized by $P/r_{h}$, but it can take a value in $10<P/r_{h}<$ 1 in the $O_{2}$ theory. At $P/r_{h}=0$, the result does not depend on $\tau$ which is located at $(\frac{k}{q\mu},\frac{T_{c}}{q\mu})\approx(0.41,0.049)$. For large enough $\tau$, $T_{c}$ linearly depends on $k$ but the coefficient is different from Eq. (32). For $\tau=1/\sqrt{3}$, we numerically obtain $T_{c}\approx 0.0193k$ for large $k$. Such behavior in $T<k/(2\sqrt{2}\pi)$ is qualitatively same as those studied in Ref. Homes' law .
@@
Finally, we show the correct phase boundaries for each theory. Combining the results in Fig. 1 for the $O_{1}$ theory, and Fig. 2 for the $O_{2}$ theory, we obtain the phase boundaries as Fig. 3. According to the critical value of $T=k/(2\sqrt{2}\pi)$, which is shown as the dotted line, we have switched the using background geometry and the corresponding results. Note that the purple curves in $T>k/(2\sqrt{2}\pi)$ are independent of the choice of $\tau$ because $\tau$ represents a coupling with the dilaton. For the $O_{1}$ theory, we employed only the purple curve corresponding to the bald black hole because all curves are located in $T>k/(2\sqrt{2}\pi)$. We also provide further comparisons with the full analysis in Appendix A. The calculation of the critical temperature $T_{c}$ in this study demonstrates that employing the probe limit yields coherent and meaningful results by duly considering the phase transition dynamics within the background gravity sector.
@@
For both the $O_{1}$ theory and the $O_{2}$ theory with large $\tau$, it is revealed that $T_{c}$ increases as the parameter $k$ grows at high $k$ regimes. Specifically, for the $O_{1}$ theory, the large-$k$ behavior can be effectively approximated using Eq. (32). Although initially, this positive correlation between $T_{c}$ and $k$, which signifies the strength of momentum dissipation, may seem counter-intuitive, such a trend has been consistently observed across various holographic models incorporating $k$ Homes' law ; tomas , as previously discussed. In light of these findings, a future discourse on the validity and implications of this scaling relationship between $T_{c}$ and $k$ from the perspective of condensed matter physics would indeed be valuable and insightful. It remains to be explored whether this phenomenon aligns with physical expectations or presents novel challenges to our current understanding.
@@ 3.2 Condensation
Figure 4 shows the condensates $\expectationvalue{O_{i}}$ as functions of $T$ for $O_{i}$ theories. We set $\tau=0$ for simplicity. To obtain this figure, we have changed the using background geometry at the phase transition point $T=k/(2\sqrt{2}\pi)$. We show this point as a small circle on the curve, if it exists below the superconducting phase transition point. The condensate $\expectationvalue{O_{1}}$ grows as temperature decreases, whereas $\expectationvalue{O_{2}}$ remains finite at $T\to 0$. In particular, there is a lower bound for the possible $T$ in the $O_{1}$ theory. This behavior is same as those studied in bhs , and is a limitation in the probe limit analysis.444 In the $O_{2}$ theory with $\tau=0$, $T\to 0$ limit is possible. Since the lowest temperature is always larger than $T=k/(2\sqrt{2}\pi)$ for the $O_{1}$ theory, the results in the bald black hole, denoted as solid curves, are favored. We expect that the low-temperature divergence of $\expectationvalue{O_{1}}$ will be cured by considering beyond the probe limit, similarly to Ref. Hartnoll:2008kx . One can see that $\expectationvalue{O_{1}}$ is enhanced by increasing $k$ for fixed $\mu$. On the other hand, $\expectationvalue{O_{2}}$ is suppressed as $k$ increases. These behaviors of the condensates are consistent with the relation between $T_{c}$ and $k$ for each theory, shown as Fig. 3.